1) Find the pH of 0.200 M HF acid solution

2) Find the pH of a and [H3O+] of 0.100 M benzoic acid solution

Explain how you got answer!
Answer #1

                   .HF + H2O    <==>   H3O+ + F-
Initial . . . . . .0.2 . . . . . . . . . . . . .0 . . . . . 0
Change . . . . .-x . . . . . . . . . . . . .x . . . . . x
At Equil. . .0.2-x . . . . . . . . . . . . .x . . . . ..x

Ka = [H3O+][F-] / [HF] = (x)(x) / (0.20-x) = 6.8 x 10^-4. . .the -x term will be small compared to 0.20 so we delete it.


x^2 / 0.20 = 6.8 x 10^-4
x^2 = 1.4 x 10^-4
x = 1.2 x 10^-2 = [H3O+]

pH = -log [H3O+] = -log (1.2 x 10^-2) = 1.93

2)